determine the slant asympotte for f(x)= x^2-5x-3 / x
Answer by HyperBrain(694)

f(x)= x^2-5x-3 / x
Simply perform the division.
We get:
f(x)=(x-5)-(3/x)
We take the quotient. So, the slant/oblique asymptote is the line y=x-5